# Evaluating Diversity in Automatic Poetry Generation

### v1

@@ 5.3 Sampling
From each model class, we randomly draw 1000 generated poems. Whenever wo do a direct comparison between training and generated data (e.g. when comparing lexical diversity), we randomly draw 10 samples of size 1000 (matching the sample size) from the train set and use mean results as representatives. We deploy this strategy to mitigate the large discrepancy in size between human data and generated poems. For all LLMs, we employ consistent decoding strategies for generation: we use the default settings of the LLaMA2 chat models on Hugging Face111https://huggingface.co/spaces/huggingface-projects/llama-2-7b-chat but limit the number of newly generated tokens to 100 for the token-level models or 300 for the character-level ByGPT5 models.44todo: 4YC: @Hannes, do you have different sampling strategies for bygpt5? H: Other than the maximum length it’s the same. In Section 7, we briefly discuss the effects of various decoding strategies.
6 Experiments and Results
We first investigate structural properties of the generated poems (repetition of instances on a surface level, length distributions, rhyming), then consider lexical and semantic properties.
@@ 6 Experiments and Results
On quatrain-level, no model exhibits memorization, with all obtaining a memorization rate of 0.
Considering couplet-level memorization, 23 out of 36 models show zero memorization, while 13 models display scores between 0.05% and 0.15%. The poetry-specific models, SA and DeepSpeare, as well as the character-level ByGPT5 models, exhibit no memorization; in contrast, GPT2 and GPTNeo models show the highest rates on average (up to 0.15% for German and 0.10% for English). When comparing models of the same architecture and training methods but varying sizes, differences are found in 6 out of 14 cases. In 5 cases, larger models have 0.05%-0.10% higher absolute memorization scores than their smaller counterparts (the English GPT2con and LLaMA2con models, and the German GPT2con, GPTNeocon, LLaMA2 models); the only exception is the English GPTNeo models, where the smaller one has a 0.05% higher memorization rate. On the other hand, there is no distinct boundary between unconditioned and conditioned trained models: in the 6 cases where discrepancies in memorization rates exist, conditioned models outperform their unconditioned ones three times, and vice versa, with absolute differences of 0.05%-0.1%.
In the verse-level evaluation, the poetry-specific models perform best overall (0.4%-0.83% for German and 0.1%-0.83% for English), followed by the ByGPT5 models (0.68%-1.3% for German and 0.58%-1.23% for English). Again, GPT2 is worst for German, exhibiting memorization rates of 4.38%-8.7%, whereas, for English, GPTNeo exhibits the highest rates, ranging from 3.5%-5.6%. Concerning different model sizes, we again see that larger models memorize more than their smaller counterparts: in 9 out of 14 cases, larger models show higher memorization rates, with an average absolute increase of 0.15%. Furthermore, unlike in the couplet-level evaluation, here, each unconditioned model exhibits a strictly higher memorization rate compared to its conditioned counterpart, with an absolute increase of 1.47% on average.
In summary, with respect to memorization: (1) Among model families, the poetry-specific and character-level models are the most diverse. (2) Larger models are less diverse compared to their smaller versions. (3) Conditional training enhances model diversity.
@@
Table 5 (appendix) reports statistics on the length of poems, both human and automatically generated. The mean length of human written poems is 28 in English and 24 in German. Histogram intersection values between samples generated by the models and the human written data range from 0.61 to 0.88 in German (LLaMA2 and SA) and from 0.48 to 0.92 in English (GPTNeo and SA). While the SA models fit the distribution of the human written poems the best, the character-level ByGPT5 models also perform well consistently with histogram intersection values between 0.77 and 0.85. The poems generated by German LLaMA2 and English GPTNeo are too short and not diverse enough (in terms of standard deviation). The poetry-specific DeepSpeare models do not match the human distribution very well either, with intersection values of 0.63 and 0.57 for German and English respectively. Here too, poems lengths are too small and not diverse enough. Conditioned models seem to fit the training data better across the board, the only exceptions being German ByGPT5 and English LLaMA2. Figure 1 illustrates the length distribution of human written poems, SA and GPTNeo for English.
@@
Figure 2 (a) shows the distributions of rhyme schemes in our human training datasets for German; Figure 3 (a) shows the corresponding distribution for English. 55todo: 5SE: Most rhymes in the training data are classified as real rhymes. For both languages, less than 15% of all quatrains in training do not rhyme at all (rhyme scheme ABCD). Excluding ABCD, the top 3 dominant rhyme schemes by appearance are ABAB, AABB and ABCB for both datasets, with a total share of approximately 60% in each language. German has a higher proportion of ABAB (above 35%), while English has ABAB and AABB in roughly equal proportions (25%). In Table 6 (appendix), we report the entropy of all rhyme distributions as well the distance between the human distribution and model distributions, measured in cross-entropy and KL divergence. The best, worst and an average model, in terms of cross-entropy, are shown in Figures 2 and 3.
Poetry-specific models: Figure 5 (appendix) shows the distributional plots for Deepspeare and SA. We observe that DeepSpeare has a very low ratio of ABCD, considerably lower than human poems (less than 5% for both languages). The three dominating patterns are AABB, ABAB, and ABBA which (only) partially agrees with the dominating patterns in the human data. Nonetheless, DeepSpeare has the best fit of all models in terms of KL divergence and almost the best fit in terms of cross-entropy. SA has a much worse fit and produces considerably too many ABCD patterns (close to or above 30% in both languages). It has one of the worst fits to the human rhyme distributions across all models.
@@
Style-conditioned LLMs are shown in Figures 8 and 9 (appendix). In general, this model class matches the human distribution closest in terms of KL divergence and cross-entropy. LLAMA3 has a particularly good fit in English, but it is noteworthy that it produces too many ABAB rhyme schemes, which may indicate a skewness due to its pre-training data. No model produces a lot of AABB rhyme pattern which abound in our human training data. Across all models in this class, the fit to the human data is still mediocre at best.
Overall, most models have clearly higher ABCD rhyming schemes than the human data, thus are underdiverse concerning rhyming. The best model class are style-conditioned LLMs (especially LLAMA3), however the poetry-specific DeepSpeare model can be considered the best individual model in terms of matching the human rhyme distribution. The character-level ByGPT5 models perform worse than word-level models without style-conditioning, but with style-conditioning they outperform the word-level models in terms of match with human rhyme distribution.
@@
Table 2 shows the lexical diversity results for English and German. For local diversity (ATTR), most of the models are close to the diversity in human-written poems, with the traditional models (DS, SA) and the Llama exceeding the ATTR values of human-written poems. For English, the least locally diverse poems are generated by GPT2-small, in the un/conditioned case respectively. For German, the least locally diverse models is GPT2neo-small, in the un/conditioned case respectively. The global diversity metrics (MATTR, MTLD) show different trends than ATTR, though. The MATTR metric suggests that most models do not generally achieve the level of diversity found in human poems: in English, only SA matches and slightly exceeds human diversity, in German, only the conditioned Llama3-large model exceeds human diversity. According to the MTLD metric, all models generate severely under-diverse output at the sample level. Here, the best model in English and German is , but even does not come close to the human level of global diversity. According to MTLD, style-conditioned LLMs consistently outperform their non-conditioned counterparts, with the English Llama2 models being the only exceptions here. ††todo: include barplots in the appendix? Moreover, we observe that model size affects all three lexical diversity metrics, whereby larger models are more diverse than their smaller counterparts. ††todo: produce a barplot that shows the effect of size? The effect of size is most pronounced for GPT2 in both languages, where ATTR, MATTR and MTLD substantially improve from the small to the larger model variant. It also noteworthy, though, that the more classical models, and , generally perform on par with recent transformer models and sometimes even outperform them, as in the case of for global diversity. This shows that unconditional LLMs avoid repetitions at a local level whereas, at the sample level, they generate poems that are lexically much more similar to each other than poems within the human sample. ††todo: explain the differences between MATTR and MTLD
@@
LLaMA3 | 59.6 / 58.2 | 68 / 62.3 |
Table 3 presents results for the semantic (cosine) similarity of quatrains: (i) within human and model-generated samples, and (ii) across generated samples and the human data. These results generally confirm the trends we found for global lexical diversity discussed above. None of the models generates a sample of poems with a within-sample diversity as low as the human with-sample diversity. SA is the model that achieves the lowest within-sample similarity and the lowest across-sample similarity, suggesting that it deviates most from the patterns in the human training data. Note that SA also achieved the best results in global lexical diversity (MATTR and MTLD in Table 2).††todo: SZ: I am not entirely sure about the across-sample similarities: is it necessarily a good thing if the models have low similarity with the human sample?. Moreover, the results on semantic similarity confirm the trends we observed with model size for lexical diversity, but disconfirm the trends for the effect of conditioning. Thus, we do not see a consistent trend for conditioned models generating samples with lower similarity/higher diversity. For model size, on the other hand, we observe a general trend towards larger models outperforming their smaller counterparts.
7 Discussion
Sampling/Searching
We deploy various decoding strategies to determine to what extent these can alter the various aspects of diversity in the generated poems. We use different combinations of temperature-based sampling (ackley1985learning), Nucleus sampling (Top-p) (holtzman2019curious) and Top-k sampling (fan-etal-2018-hierarchical) as sampling strategies and further deploy two variants of contrastive search (su2022contrastive). Results indicate that the various techniques can only slightly increase diversity in one or more aspects. Moreover, aggressive sampling often leads to output degeneration, causing the models to (partially) repeat verses in a quatrain. We provide some examples in the appendix, see Tables 7 to 12 as well as Figures 10 to 15.
@@ 7 Discussion
We have seen that unconditioned LLMs exhibit poor results with regard to different dimensions of diversity: they do not rhyme, are lexically underdiverse and do not show sufficient semantic variation. However, character-level models are more diverse than word level models. Style-conditioned models perform better regarding rhyming, semantic variation, and lexical variation but word level style-conditioned models are prone to severe memorization from the training data, in particular when the model is large and the training set is small. Character-level style-conditioned LLMs produce overall best diversity results and do not deteriorate as a function of model/training data size. In terms of diversity, poetry-specific Deepspeare performs similar as character-level LLMs but requires more modeling effort from human experts (e.g., in developing rhyming components).
@@ 8 Conclusion
To date, evaluation of automatic poetry generation has almost exclusively focused on human evaluation and forms of the Turing test. Our work shows that an automatic assessment of the diversity of generated poems covers an important blind spot of existing studies. Our evaluations shed light on the fact that none of the state-of-the-art poetry generators is able to match the level of diversity in human poems, confirming previous evaluations of diversity in other NLG tasks ippolito-etal-2019-comparison; schuz-etal-2021-diversity; stasaski-hearst-2022-semantic. Our study also adds a new dimensions to previous work on diversity, by showing that diversity on the level of rhyming is particularly hard to achieve for neural generators and interacts with other dimensions of diversity in poetry generation, i.e., style conditioned LLMs do not only achieve a better match with human rhyme and length distributions, but also higher lexical and semantic diversity . We also find that memorization — a general and widely discussed limitation of LLMs (carlini2021extracting) — is a potential issue in poetry generation, especially for certain combinations of model sizes and finetuning schemes, complementing existing studies in this area mireshghallah-etal-2022-empirical.
We release all code upon acceptance.
@@
Generally, our work is concerned with the evaluation of NLG systems; evaluation methods and evaluation metrics (zhao-etal-2019-moverscore; Zhang2020BERTScore; NEURIPS2021_e4d2b6e6; chen_menli:2023; peyrard-etal-2021-better) are a well-known and notorious issue in this research field. While a lot of recent work has aimed at improving common practices in human evaluation (belz-etal-2023-missing) or advancing the study of metrics for quality or fluency of NLG outputs, the evaluation of diversity is comparatively under-researched. In this work, we aimed at providing a range of metrics assessing different aspects of diversity, but could not cover all potentially interesting ways of measuring diversity. Here, future work could look at further aspects of formal and structural diversity (e.g. at the level of syntax, or meter), or other aspects of semantic diversity (e.g. topical diversity, rhetorical figures). Future work could also consider more (diverse) languages and other genres and datasets for poetry.
@@
Often, the discussion of creative AI systems in public discourse is surrounded by misconceptions, hypes and even myths (veale2012exploding). Our work contributes to a careful operationalization and objective assessment of the creative capbalities of AI systems in the area of poetry generation.
All the datasets, models and code used in this work are publicly available or will be made available upon publication. We have not collected private or sensitive data and have only used language models with free access, such that our experiments can be fully replicated by anyone.

### v2

@@ 5.1 Diversity Evaluation
From each model, we randomly draw 1000 generated poems. Whenever we do a direct comparison between training and generated data (e.g. when comparing lexical diversity), we randomly draw 10 samples of size 1000 (matching the sample size) from the train set and use mean results as representatives. We deploy this strategy to mitigate the large discrepancy in size between human data and generated poems.
We first investigate structural properties of the generated poems (repetition of instances on a surface level, length distributions, rhyming), then consider lexical and semantic properties. After discussing each dimension of diversity, we provide a brief summary that generalizes across different model classes (e.g., poetry-specific vs. style conditioned vs. unconditioned, character- vs. word-level, larger vs. smaller). These summaries are based on Table 3.
@@
| DE | EN | |||
| verse | couplet | verse | couplet | |
| DeepSpeare | 0.83% | 0.83% | ||
| SA | 0.40% | 0.10% | ||
| ByGPT5 | 1.30%∗ | 1.23%∗ | ||
| ByGPT5 | 1.23% | 0.93% | ||
| GPT2 | 6.85% | 0.10% | 3.90% | 0.10% |
| GPT2 | 8.70%∗ | 0.10% | 4.03%∗ | 0.10% |
| GPTNeo | - | 5.60%∗ | 0.05% | |
| GPTNeo | - | 4.73% | 0.10%∗ | |
| LLaMA2 | 4.65% | 3.45%∗ | 0.05%∗ | |
| LLaMA2 | 5.45%∗ | 2.48% | ||
| LLaMA3 | 3.60% | 2.88% | 0.05% | |
| ByGPT5 | 0.90%∗ | 0.58% | ||
| ByGPT5 | 0.68% | 0.75%∗ | ||
| GPT2 | 4.38% | 0.15%∗ | 2.33%∗ | 0.10%∗ |
| GPT2 | 6.90%∗ | 0.10% | 2.03% | |
| GPTNeo | - | 3.88%∗ | 0.05%∗ | |
| GPTNeo | - | 3.50% | ||
| LLaMA2 | 4.03%∗ | 0.05%∗ | 2.23%∗ | |
| LLaMA2 | 0.70% | 0.55% | ||
| LLaMA3 | 2.33% | 1.65% | ||
Table 4 showcases the couplet- and verse level memorization rates. Since all models exhibit zero memorization rates on quatrain-level, we omit them in the table.
Considering couplet-level memorization, 23 out of 36 models show zero memorization, while 13 models display scores between 0.05% and 0.15%. The poetry-specific models, SA and DeepSpeare, as well as the character-level ByGPT5 models, exhibit no memorization; in contrast, GPT2 and GPTNeo models show the highest rates on average (up to 0.15% for German and 0.10% for English). When comparing models of the same architecture and training methods but varying sizes, differences are found in 6 out of 14 cases. In 5 cases, larger models have 0.05%-0.10% higher absolute memorization scores than their smaller counterparts (the German GPT2con and LLaMA2con models, and the English GPT2con, GPTNeocon, LLaMA2 models); the only exception is the English GPTNeo models, where the smaller one has a 0.05% higher memorization rate. On the other hand, conditioned models mostly outperform their unconditioned counterparts: in 4 out of 6 cases where discrepancies in memorization rates exist, the conditioned ones exhibit lower memorization rates, with absolute declines of 0.05%-0.10%.
In the verse-level evaluation, the poetry-specific models perform best overall (0.4%-0.83% for German and 0.1%-0.83% for English), followed by the ByGPT5 models (0.68%-1.3% for German and 0.58%-1.23% for English). SA is the best individual model, obtaining memorization rates of 0.4% for German and 0.1% for English. Again, GPT2 is worst for German, exhibiting memorization rates of 4.38%-8.7%, whereas, for English, GPTNeo exhibits the highest rates, ranging from 3.5%-5.6%. Concerning different model sizes, we again see that larger models memorize more than their smaller counterparts: in 9 out of 14 cases, larger models show higher memorization rates, with an average absolute increase of 0.15%. Here, each conditioned model exhibits a strictly lower memorization rate compared to its unconditioned counterpart, with an absolute decrease of 1.47% on average.
Overall: (1) No models exhibit severe memorization issues, such as copying entire poems or large portions of poem snippets from the training data. In terms of memorization, (2) among model groups, the poetry-specific and character-level models are more diverse; SA is the best individual one. (3) Larger models are less diverse compared to their smaller versions. (4) Conditional training enhances model diversity.
@@
Table 7 (appendix) reports statistics on the length of poems, both human and automatically generated. The mean length of human written poems is 28 in English and 24 in German. Histogram intersection values between samples generated by the models and the human written data range from 0.61 to 0.88 in German (LLaMA2 and SA) and from 0.48 to 0.92 in English (GPTNeo and SA). While the SA models fit the distribution of the human written poems the best, the character-level ByGPT5 models also perform well consistently with histogram intersection values between 0.77 and 0.85. The poems generated by German LLaMA2 and English GPTNeo are too short and not diverse enough (in terms of standard deviation). The poetry-specific DeepSpeare models do not match the human distribution very well either, with intersection values of 0.63 and 0.57 for German and English, respectively. Here, too, poem lengths are too short and not diverse enough. Conditioned models seem to fit the training data better across the board, the only exceptions being German ByGPT5 and English LLaMA2. Figure 4 (appendix) illustrates the length distribution of human written poems, SA and GPTNeo for English.
Overall, regarding the alignment with human distributions: (1) Character-level ByGPT5 models generally align best with human data, followed by poetry-specific models; nevertheless, the poetry-specific SA is the top individual model. (2) Style-conditional models outperform the unconditioned trained ones. (3) Smaller models demonstrate a better fit than the larger ones.
@@
Figures 2 (a) and 3 (a) show the distributions of rhyme schemes in our human training datasets for German and English, respectively. For both languages, less than 15% of all quatrains in training do not rhyme at all (rhyme scheme ABCD). Excluding ABCD, the top 3 dominant rhyme schemes by appearance are ABAB, AABB and ABCB for both datasets, with a total share of approximately 60% in each language. German has a higher proportion of ABAB (above 35%), while English has ABAB and AABB in roughly equal proportions (25%). Table 8 (appendix) reports the entropy of all rhyme distributions and the distance between the human distribution and model distributions, measured in KL divergence. The best, worst and an average model, in terms of KL divergence, are shown in Figures 2 and 3.
Poetry-specific models: Figure 5 (appendix) shows the distributional plots for DeepSpeare and SA. We observe that DeepSpeare has a very low ratio of ABCD, considerably lower than human poems (less than 5% for both languages). The three dominating patterns are AABB, ABAB, and ABBA which (only) partially agrees with the dominating patterns in the human data. Nonetheless, DeepSpeare has the best fit of all models in terms of KL divergence, ranking first for German and second for English. SA has a much worse fit and produces considerably too many ABCD patterns (close to or above 30% in both languages). It has one of the worst fits to the human rhyme distributions across all models.
@@
Style-conditioned LLMs are shown in Figures 8 and 9 (appendix). In general, this model class matches the human distribution closest in terms of KL divergence. However, no model produces a lot of AABB rhyme pattern which abound in our human training data. Across all models in this class, the fit to the human data is still mediocre at best.
Overall, most models have clearly higher ABCD rhyming schemes than the human data, thus are underdiverse concerning rhyming. (1) Conditioned models very clearly outperform unconditioned models and (2) character-level and poetry-specifc models are clearly better than word-level models in terms of matching the human rhyme distribution. (3) There is no clear size effect.
@@
Table 5 shows the lexical diversity results for English and German. For local diversity (ATTR), most of the models are close to the diversity in human-written poems, with the traditional models (DeepSpeare, SA) and the LLaMA exceeding the ATTR values of human-written poems. For German, the least locally diverse poems are generated by GPT2, in the un/conditioned case, respectively. For English, the least locally diverse models is GPTNeo, in the un/conditioned case, respectively. The global diversity metrics (MATTR, MTLD) show different trends than ATTR, though. The MATTR metric suggests that most models do not generally achieve the level of diversity found in human poems: in English, only SA matches and slightly exceeds human diversity, in German, only the LLaMA2 and LLaMA3 model exceeds human diversity. According to the MTLD metric, all models generate severely under-diverse output at the sample level. Here, the best model in English and German is SA, but even SA does not come close to the human level of global diversity. According to MTLD, style-conditioned LLMs consistently outperform their non-conditioned counterparts, with the English LlaMA2 models being the only exceptions here. Moreover, we observe that model size affects all three lexical diversity metrics, whereby larger models are more diverse than their smaller counterparts. The effect of size is most pronounced for GPT2, where ATTR, MATTR and MTLD substantially improve from the small to the larger model variant. Generally, the MTLD results suggest more pronounced differences between models as well as humans and models than MATTR.
Overall, in terms of lexical diversity, (1) neural models match human performance at the local level but fall short at the global level. (2) Poetry-specific models outperform other model classes, while character-level LLMs are most deficient (except for MTLD). (3) Conditional training is beneficial. (4) Larger models perform better.
@@
| LLaMA3 | 59.6 / 58.2 | 68.0 / 62.3 |
Table 6 presents results for the semantic (cosine) similarity of quatrains: (i) within human and model-generated samples, and (ii) across generated samples and the human data. None of the models generates a sample of poems with a within-sample diversity as low as the human with-sample diversity. SA is the model that achieves the lowest within-sample similarity and the lowest across-sample similarity.
Overall, (1) poetry-specific models are most diverse regarding semantic similarity and word-level models are least diverse; (2) style-conditioning makes models slightly more diverse semantically; (3) larger models are also slightly more diverse.
@@
We have seen that unconditioned LLMs exhibit poor results across various dimensions of diversity: they often do not rhyme, are lexically underdiverse and do not show sufficient semantic variation. However, character-level models are more diverse than word level models. Style-conditioned models perform better regarding memorization, rhyming, and lexical variation, while deviating less from human poems according to the distribution match of length and rhymes. On the other hand, larger LLMs often outperform their smaller counterparts in semantic and lexical diversity, but they also tend to memorize more from the training data. Character-level style-conditioned LLMs produce overall best diversity results and do not deteriorate as a function of model/training data size. In Appendix A.3, we calculate the average ranks of the models across all 5 dimensions, finding that indeed, for both languages, the conditioned trained ByGPT5 models perform overall best among all models, ranking as the first and second places for German and the first and third places for English. In terms of diversity, poetry-specific SA and DeepSpeare overall lag only slightly behind character-level LLMs but require more modeling effort from human experts (e.g., in developing rhyming components). The largest word-level LLMs explored in this work, LLaMA2 and LLaMA3, generally perform best among the word-level models; however, they do not exhibit superiority over the style-conditioned character-level models and poetry-specific models as well.
We also compute Pearson’s correlations between ranks for different dimensions. For German, the highest correlation is between semantic diversity and memorization (0.842), followed by the two moderate to high correlations: 0.526 (semantic vs. lexical) and 0.518 (memorization vs. rhyme). Two pairs show moderate correlations: 0.480 (semantics vs. length) and 0.404 (memorization vs. rhyme). The remaining pairs exhibit weak positive or negative correlations, with absolute values between 0.051 and 0.228. For English, no pairs exhibit high correlations. Two pairs show moderate to high correlations: 0.628 and 0.635 (memorization vs. semantics/length). Three pairs demonstrate moderate correlations, ranging from 0.307 to 0.357 (semantics vs. lexical/length and memorization vs. length). The others show weak correlations, with absolute values between 0.024 and 0.267. Concretely, these sometimes low correlations are mirrored in the different ranks models have across different dimensions: for example, SA is almost as diverse as the human training data regarding semantics and length, but provides one of the worst fits regarding rhyming. This indicates that most current models face a tradeoff for different diversity dimensions.
5.2 Quality Evaluation
Diversity in model outputs could sometimes result from low coherence or a lack of meaningful content. To investigate whether this is the case, we conducted a small-scale human evaluation of the overall quality of quatrains, focusing specifically on coherence and semantics (punctuation was omitted here, as it was also excluded during the diversity evaluation). In this evaluation, we compared 60 outputs across 5 systems (12 outputs per system) for each language, including human-written quatrains, and the outputs of the winning models in overall, lexical, semantic, and rhyme diversity (as presented in Tables 9 and 10 in the appendix). We created 15 annotation instances; in each instance, an annotator was given 4 quatrains and asked to select both the best and the worst among them.
As Table 11 in the appendix displays, for German, human quatrains are clearly preferred (they were chosen as the best 12 times and the worst 0 times). The best automatic system is the overall winning ByGPT5 model (best 2 times; worst 1 time); SA is the worst (worst 8 times). For English, the lexical winning LLaMA3 model is the best in terms of coherence (best 6 times; worst 0 times), followed by the rhyme winning GPTNEO model (best 5 times; worst 0 times); SA is again the worst (worst 11 times). However, we noted that our evaluator was a native speaker of German but not English and said that the German evaluation was much easier for him. The older SA model appears to have higher diversity at the cost of quatrain quality. However, overall, we conclude that more diverse models also seem to be qualitatively better — this does not have to be a causal/strong relationship, however, especially for the newer LLMs. Tables 12 and 13 in the appendix present 10 sample quatrains selected as the best in our human evaluation, including both system-generated and human-written ones.
@@ 6 Conclusion
Our work is the first and most comprehensive automatic evaluation of poetry diversity, yielding several interesting observations: for example, we find that style-conditioning enhances virtually all measures of diversity and that character-level modeling also increases diversity, including reducing memorization. Our evaluations also shed light on the fact that none of the state-of-the-art poetry generators is able to match the level of diversity in human poems. Thus, we find overall that an automatic assessment of the diversity of generated poems covers an important blind spot of existing studies. Future work should aim for more diverse automatic poetry generation systems as a prerequisite of general computational creativity.
@@
Due to the hardware constraints and time limitations, we did not run experiments multiple times to take the averages or optimize the training hyperparameters, which may have introduced a degree of randomness in our results. For example, in our initial experiments, we trained GPT2 models with a slightly different setting. Compared to the GPT2 models we mainly reported, these models behave slightly differently. E.g., they exhibit better lexical diversity, as shown by an increase in ATTR from 0.87 to 0.89, MATTR from 0.84 to 0.86, and MTLD from 88 to 101 on average. Similarly, they are also more diverse according to the semantic similarity metrics, which are on average 0.02-0.03 lower. In contrast, these models perform worse in rhyming; they have a 10% lower chance of producing rhymed quatrains, and their rhyme distributions are more distant from human distributions (0.27 higher KL divergence). Despite these differences, our findings are generally robust as we report averages over model classes in our analysis. For the same reason, we did not select the largest versions of these models; nevertheless, our evaluation already shows prominent differences in diversity across model sizes.
Further, we note that our trained LLMs occasionally do not generate texts in the form of a quatrain (i.e., 4 verses). These outputs were excluded from the analysis, though such cases are rare (1.5% on average).
@@
All the datasets, models, and code used in this work will be made publicly available. We have not collected private or sensitive data and have only used language models with free access, such that our experiments can be fully replicated by anyone.
Generally, our work is concerned with the evaluation of NLG systems; evaluation methods and evaluation metrics (Zhao et al., 2019; Zhang et al., 2020; Peyrard et al., 2021; Yuan et al., 2021; Chen et al., 2022; Chen and Eger, 2023; Leiter et al., 2023) are a well-known and notorious issue in this research field. While a lot of recent work has aimed at improving common practices in human evaluation (Belz et al., 2023) or advancing the study of metrics for quality or fluency of NLG outputs, the evaluation of diversity is comparatively under-researched. In this work, we aimed at providing a range of metrics assessing different aspects of diversity, but could not cover all potentially interesting ways of measuring diversity. Here, future work could look at further aspects of formal and structural diversity (e.g. at the level of syntax, or meter), or other aspects of semantic diversity (e.g. topical diversity, rhetorical figures). Future work could also consider more (diverse) languages and other genres and datasets for poetry.